# MetaML-Pro: Cross-Stage Design Flow Automation for Efficient Deep Learning Acceleration

### v2

@@ 4.6. Design Space Exploration (DSE)
Within our approach, Design Space Exploration (DSE) algorithms are critical for automatically adjusting the optimization parameters in subsequent iterations to improve design performance, while maintaining high model accuracy and meeting the constraints of the device.
Bayesian optimization is a key algorithm we utilize for global optimization of expensive-to-evaluate functions (represented as f(x)). It models the objective function using a probabilistic model (e.g., Gaussian Process) and uses an acquisition function to decide where to sample next, balancing exploration of the design space with exploitation of known optimized configurations. In our context, x denotes the array of configuration settings, which serve as parameters for each optimization task within our framework. The process guides towards the most promising configurations for evaluation until a predetermined number of iterations is reached.
Our performance metrics encompass accuracy, hardware resource utilization, and adherence to operational constraints. We leverage hardware usage data obtained through Xilinx Vivado, including latency, DSP, LUT, FF, and BRAM, to inform our optimization strategy for model deployment, adjusting the optimization parameters including scaling, pruning, and quantization thresholds ($\alpha_{p/s/q}$). Due to the heterogeneity of these metrics, direct summation is impractical, necessitating normalization to standardize them. This normalization technique is widely applied in various DSE methods (Sohrabizadeh et al., 2022).
Moreover, when optimizing a complex system, such as FPGA-based DNN accelerators, practical constraints must be considered. These constraints may arise from hardware limitations, performance requirements, and other critical factors. Incorporating these constraints into the Bayesian optimization process ensures that our search for the optimal solution occurs within a feasible and practical region of the design space. For hardware utilization constraints, a score becomes a negative maximum integer when the utilization is higher than a predefined threshold. This signals the Bayesian algorithm that the input parameter is unsuitable, steering the optimization away from poor configurations. The same approach applies to model accuracy constraint. The pseudocode of Bayesian score algorithm is shown below:
if (Bayesian constraints not met) then
$f(x)=-sys.maxsize$
else
$f(x)=\sum_{x\in\{Acc.,DSP,LUT,BRAM\}}Norm\_Results[x]\times W[x]$
The algorithm assigns weights (W[metric]) to each metric to indicate their relative importance.
@@
LSTM (Que et al., 2020) MNIST Sequence classification
@@ 5.1. Experimental Setup
Experiments were conducted in Python 3.9.15 with benchmark workloads from typical DNN applications, as presented in Table 2), including jet identification (Duarte et al., 2018; Moreno et al., 2020), image classification using VGG(Simonyan and Zisserman, 2014) and ResNet (He et al., 2016) networks. The jet identification task targeted FPGA-based CERN Large Hadron Collider (LHC) triggers with a 40 MHz input rate and a response latency of less than 1 microsecond. Default frequencies were 100MHz for Zynq 7020 and 200MHz for Alveo U250 and VU9P, and the HLS4ML task used 18-bit fixed-point precision with 8 integer bits.
@@ 5.2. Single $O$-task Strategies
Figure 9. The auto-pruning algorithm applied to models with binary search direction shown. Omitting step s1 for visibility. The blue arrow indicates an accuracy loss $>$ user threshold; red denotes the optimal pruning rate.
@@
Figure 11. (a) Quantized bitwidth of each layers in VGG7. (b) Quantized bitwidth of each layers in JetDNN.
@@
The effectiveness of the auto-pruning algorithm is demonstrated in Fig. 10. Figures (a) and (c) depict the pruning rate and accuracy for Jet-DNN and ResNet9 in each step, while (b) and (d) show the resources utilization on Zynq 7020 and U250. As the pruning rate increases, hardware resource requirements, particularly DSPs and LUTs, decrease, leading to improved FPGA performance. The design candidate with the highest pruning rate within the allowed tolerance is selected.
@@
This section showcases the evaluation results of the quantization (Q) optimization applied to multiple DNN models within the hardware (HLS) optimization space.
Fig 11 shows the precision of the weights, biases and output of each virtual layer of the VGG7 model after being tuned by the quantization strategy with $\alpha_{q}$ set to 1%. Table 3 shows how the quantization affected key evaluation metrics relating to the performance and resource usage of 2 DNN designs. With $\alpha_{q}$ set to 1%, the proposed QHS quantization algorithm reduces DSP usage by a factor of 4.9 for the VGG7 model, and 8.5 for the JetDNN model. When $\alpha_{q}$ is increased to 5%, the designs are further compacted, but with a larger real accuracy loss. This table highlights how varying $\alpha_{q}$ levels affect model accuracy, DSP, LUT, and FF usage, illustrating the trade-offs between resource savings and accuracy as well as the effectiveness of the proposed QHS quantization algorithm.
@@ 5.3. Multiple $O$-task Strategies
With our framework, new strategies can be derived by building and revising a design-flow architecture. For instance, by inserting a scaling $O$-task before the pruning $O$-task in Fig. 8(a), a custom combined strategy can be created with results shown in Fig. 10(e) and (f). The new optimal pruning rate is 84.4% (Fig. 8(e)), lower than the previous 93.8% (Fig. 8(a)), due to reduced redundancy from the preceding scaling task. By switching the order of the $O$-tasks, a different optimization strategy performing pruning-then-scaling (pruning $\rightarrow$ scaling) is achieved, resulting in a 0.7% accuracy drop after one scaling step, as seen in Fig. 10(g). Moreover, the three optimization $O$-tasks, pruning, scaling, and quantization, can be integrated into a single automated cross-stage strategy to enhance both performance and hardware efficiency, as illustrated in Fig. 8(c). We discuss the effects of different design-flow architectures with various combinations and orders in Section 5.7, and various tolerable loss in Section 5.8
@@ 5.4. Branching Flow
Fig. 13 presents the results of applying this design flow to an LSTM model on the MNIST dataset. Specifically, Fig. 13(a) illustrates the pruning rate and accuracy at each step using the PRUNING $O$-task. The tolerance is set to less than $\alpha_{p}$ (2%) in this design. Figs 13(b) and (c) show the resource utilization of the LSTM design after each pruning step respectively on an AMD KU115 FPGA and on an Intel A10 1150 FPGA. The DSP consumption is reduced from 6011 (108%) to 2101 (38.1%) on the AMD FPGA after the final pruning rate is optimized to be 71.9%. Compared to AMD’s HLS compiler, which prefers DSP blocks, Intel’s HLS compiler tends to favor the use of soft multipliers for implementation. As shown in Fig 13(c), most of the computation kernels are implemented using logic resources rather than DSP blocks.
While this design flow currently supports two types of FPGAs, it can be extended to include additional paths such as GPU, CPU and ASIC technologies. Moreover, this evaluation underscores the flexibility of our approach in utilizing the same software optimization task, specifically PRUNING, across multiple hardware targets.
@@ 5.6. Bottom-up Flow
Fig 15 reveal two distinct flows in the optimization strategies: top-down, where information flows from the DNN to the hardware stage, and bottom-up, the reverse. We have automated both flows by customizing the $K$-task BRANCH with a user-defined predicate function to activate or stop the bottom-up flow if the resulting design overmaps. The BRANCH task also support a user-supplied action function that is triggered when the predicate condition is true. For our current strategies, the action function changes the CFG section of the meta-model, increasing the accuracy tolerance parameters $\alpha_{p}$, $\alpha_{s}$, and $\alpha_{q}$ (if applicable) for the next iteration. Note that our current strategies in Fig. 15 have two DSE loops: an inner-loop that is codified within each $O$-task and an outer-loop supported by the bottom-up flow. Users can develop more complex strategies by customizing the outer loop, for instance, by updating the bottom-up condition and parameter tuning.
@@ 5.7. Combination and Order
The search for optimal combinations and sequences with multiple $O$-tasks can be computationally demanding due to the exponential growth of the search space with the number of $O$-tasks. Advanced search methods can replace the current brute-force search. Recent studies (Kurek et al., 2016; Ferretti et al., 2022; Wu et al., 2021) demonstrate surrogate models can expedite the search. Our framework facilitates the deployment of such techniques, enabling users to interface an search algorithm with the hardware design with minimal programming effort.

### v3

@@ 4.6. Design Space Exploration (DSE)
1
Input: Optimization tasks $O=\{\text{SCALING (S)},\text{PRUNING (P)},\text{QUANTIZATION (Q)}\}$;
Orderings $\Pi\subseteq\text{Perm}(O)$ (e.g., $\{S{\to}P{\to}Q,\;P{\to}S{\to}Q,\dots\}$);
Initial parameters $\theta_{1}=\tau_{1}=(\alpha_{p},\alpha_{s},\alpha_{q})$;
Resource/latency/accuracy limits $u_{\max}$, $t_{\max}$ $acc\_loss_{\max}$;
BO budget $T$; initial design set $\mathcal{D}_{0}$; vendor set $\mathcal{V}=\{\text{AMD},\text{Intel}\}$;
Meta-model MMMM with sections CFG/LOG/MODEL-SPACE.
Output: Best design(s) $\mathcal{D}^{\star}$ and Pareto set $\mathcal{P}$.
2
31exInitialize: Fit surrogate $\mathcal{S}$ on $\mathcal{D}_{0}$; Set BO results $\mathcal{E}_{\pi}\leftarrow\emptyset$ for all $\pi\in\Pi$.
4 for $\pi\in\Pi$ in parallel do // FORK over orderings
5  for $t\leftarrow 1$ to $T$ do
6     $m_{t}\leftarrow\textsf{Evaluate}(\pi,\theta_{t},MM)$
7     if $\textsf{Feasible}(m_{t},\theta_{t},u_{\max},t_{\max},acc\_loss_{\max})$ then
8        $s_{t}\leftarrow\textsf{Score}(\textsf{Normalize}(m_{t}))$
9
10    else
        $s_{t}\leftarrow-M$
        // constraint-violated candidate receives large negative score
11
    $\mathcal{S}\leftarrow\textsf{UpdateSurrogate}(\mathcal{S},\theta_{t},s_{t})$
     // Fit surrogate
12     $\mathcal{E}_{\pi}\leftarrow\mathcal{E}_{\pi}\cup\{(\theta_{t},m_{t})\}$
     $\theta_{t+1}\leftarrow\textsf{ProposeCandidate}(\mathcal{S},\text{CFG}(MM),\pi)$
     // acquisition-guided
13
14
$\mathcal{P}\leftarrow\textsf{ParetoReduce}\!\left(\bigcup_{\pi\in\Pi}\mathcal{E}_{\pi}\right)$
// REDUCE across paths
15 $\mathcal{D}^{\star}\leftarrow\textsf{SelectBest}(\mathcal{P},\text{policy})$
16 return $(\mathcal{D}^{\star},\mathcal{P})$
17
[…]
  // Top-down pass with inner DSE inside each $O$-task
19  $\textsf{Apply}(MM.\text{CFG},\theta)$
20  $\textsf{Run}(\text{KERAS-MODEL-GEN})$;
21  foreach $O\in\pi$ do
     $\textsf{Run}(O)$
     // SCALING/PRUNING/QUANT.: each has its own inner DSE
22
23 $\textsf{Run}(\text{HLS4ML})$
  // Vendor-aware BRANCH
24  if $\textsf{VendorSelect}(MM)=\text{AMD}$ then
25     $\textsf{Run}(\text{VIVADO-HLS})$
26 else
27     $\textsf{Run}(\text{INTEL-oneAPI})$
28 $m\leftarrow\textsf{CollectMetrics}(\text{accuracy},\text{latency},\text{DSP},\text{LUT},\text{FF},\text{BRAM})$
29  return $m$
30
[…]
32  return $\big(m.\text{util}\leq u_{\max}\big)\;\textnormal{{and}}\;\big(m.\text{lal}\leq t_{\max}\big)\;\textnormal{{and}}\;\big(m.\text{acc\_loss}\leq acc\_loss_{\max}\big)$
33
34Function Score(m~)($\tilde${m}))
35  return $(w_{1}\,\tilde{m}.\text{Acc}\;+\;w_{2}\,\tilde{m}.\text{DSP}\;+\;w_{3}\,\tilde{m}.\text{LUT}\;+\;w_{4}\,\tilde{m}.\text{lat})$
36
Algorithm 1 MetaML-Pro: Cross-Stage Constrained BO with Bottom-Up Feedback and Order Search
@@ 27     $\textsf{Run}(\text{INTEL-oneAPI})$
Within our approach, Design Space Exploration (DSE) algorithms are critical for automatically adjusting the optimization parameters in subsequent iterations to improve design performance, while maintaining high model accuracy and meeting the constraints of the device. We use Bayesian optimization to select high-level configuration parameters $\boldsymbol{\theta}$ for the cross-stage flow, e.g., global tolerances $\boldsymbol{\tau}=(\alpha_{p},\alpha_{s},\alpha_{q})$. Layer-wise details are decided inside each $O$-task’s inner search (Sections 4.2–4.4), while BO governs the outer loop (Algorithm 1).
4.6.1. Surrogate model.
We model the scalarized objective $f(\boldsymbol{\theta})$ with a zero-mean Gaussian Process (GP) using a Matérn covariance function and homoscedastic Gaussian noise. Inputs are min-max normalized to $[0,1]$. GP hyperparameters are re-estimated by maximizing the marginal likelihood at each iteration using the current design set.
4.6.2. Objective scalarization and normalization.
Let results be $\text{Acc},\text{DSP},\text{LUT},\text{Lat}$. We normalize each metric to $[0,1]$ using ranges observed so far, then form the minimization objective
[…]
subject to feasibility. Unless otherwise stated we use equal weights. To target particular frontiers we choose weights accordingly; e.g., Accuracy-DSP uses $(w_{1},w_{2},w_{3},w_{4})=(0.5,0.5,0,0)$, Accuracy-LUT uses $(0.5,0,0.5,0)$, and the balanced Acc-DSP-LUT-Lat setting uses $(0.25,0.25,0.25,0.25)$.
4.6.3. Constraints.
Feasibility requires: (i) device utilization within limits (DSP/LUT), (ii) latency $\leq t_{\max}$, and (iii) model accuracy loss $\leq Acc\_loss_{max}$. We treat constraints as hard: infeasible candidates are rejected and scored as M-M in Algorithm 1’s Score function.
4.6.4. Acquisition.
The Bayesian optimization algorithm finds next candidate in the search space by maximizing the acquisition function’s value, where the acquisition function is Expected Improvement (EI).
4.6.5. Convergence.
The outer BO loop stops when any of the following holds: (i) budget $T=22$ iterations per ordering $\pi$, or (ii) the incumbent design remains unchanged for 5 iterations. Inner loops terminate as defined earlier: pruning uses a binary search until $|\Delta\text{rate}|<\beta_{p}$ (Section 4.2), scaling stops when loss exceeds $\alpha_{s}$ (Section 4.3), and QHS halts when no further reducible precisions remain under $\alpha_{q}$ (Section 4.4).
4.6.6. Global tolerance variables.
We treat the vector of global tolerances $\tau=(\alpha_{p},\alpha_{s},\alpha_{q})$ as part of the BO decision variable $\boldsymbol{\theta}$. The user supplies only an initial $\tau_{0}$ and bounds $\tau_{\min},\tau_{\max}]$. At each BO iteration, the acquisition proposes a new $\boldsymbol{\theta}_{t}$ that includes $\tau_{t}$; no per-benchmark manual tuning of $\alpha_{x}$ is required. Inner searches remain local to each $O$-task, while BO governs the outer loop over $\tau$ and flow-level choices (Algorithm 1).
4.7. DSE-Generated Hardware Architecture
We adopt a dataflow-style, layer-pipelined streaming hardware architecture, which is well-suited for implementing DNN designs. All candidates explored by MetaML-Pro are instantiated from the hls4ml hardware library as per-layer compute kernels (e.g., nnet::conv_2d_cl, and nnet::dense) connected by streaming channels. Concurrency (e.g., unrolling/reuse, initiation interval) and buffering (e.g., FIFO depths) are driven by HLS pragmas and layer configuration parameters. The library is initially configured via IOType, default_precision, device, and clock (Table 1). In addition, the generated HLS can be refined by metaprogramming that inserts/tunes pragmas and C++ codes through AST manipulation (Section 4.5).
This architecture adopts a layer-wise pipeline (Nakahara et al., 2020; Que et al., 2022a, 2024b; Sun et al., 2025a), as shown in Fig. 8(b). Each major computational block of DNN models is mapped onto a tailored hardware module, with local memory connected via on-chip streams. The modules are chained into a deep pipeline so that different layers process different inputs concurrently, enabling continuous dataflow. In contrast to a conventional single-engine design (Que et al., 2020b, 2021a; Fan et al., 2021) that time-multiplexes a common compute core across all layers (Fig. 8(a)), this architecture allows each block in DNN to be independently optimized for resource utilization, numerical precision, and micro-architecture. Intermediate data remain on-chip between layers to avoid repeated off-chip memory accesses and reduce inference latency, while exploiting the full potential of FPGA customizability. Please refer to (Que, 2023) for a more detailed discussion of these two architectures.
Figure 8. FPGA-based DNN accelerator architectures: (a) A single-engine hardware architecture with a shared computation engine that is time-multiplexed across all layers; (b) A stream-based layer-wise pipeline hardware architecture with several custom engines, each processing a sub-layer or a whole layer or even multiple layers in a DNN model.
In regards to the DSE process, our proposed Bayesian-optimization (BO) loop treats the global tolerance vector $\tau=(\alpha_{p},\alpha_{s},\alpha_{q})$ as part of the decision variable and automatically proposes new $\tau_{t}$ each iteration. For each BO proposal, the $O$-tasks run inner searches that directly modify the hardware: (i) Pruning deletes weights/filters, shrinking convolutions and dense layers; (ii) Scaling reduces layer sizes; (iii) Quantization (QHS) assigns mixed-precision formats per virtual layer at the HLS C++ level. The outcome is an HLS-based C++ design with a specific level of parallelism, buffering, and custom numeric formats, ready for synthesis and evaluation.
4.8. Collaborative Optimization Mechanics
MetaML-Pro realizes collaboration across four interacting layers spanning software and HLS domains. First, it employs a data/state-level collaboration via the meta-model. All tasks (software and HLS) read/write a shared meta-model (CFG/LOG/MODEL-SPACE), which carries models, tool reports, and computed metrics across stages. This allows later-stage hardware outcomes (latency/utilization) to inform earlier software optimizations programmatically. Second, it employs control-level collaboration via $K$-tasks. BRANCH triggers bottom-up adaptation when constraints are violated; FORK/REDUCE explore alternative orderings and merge by Pareto dominance; JOIN composes multi-stage flows. This moves beyond static pipelines to runtime flow reconfiguration. Third, it has parameter-level collaboration via Bayesian Optimization based outer loop and inner local searches. The outer constrained BO jointly chooses the global tolerance vector $\tau=(\alpha_{p},\alpha_{s},\alpha_{q})$ and searches across task orderings $\Pi$; each O-task then runs its own inner DSE (binary-search pruning, progressive scaling, QHS quantization). This couples global, cross-stage choices with local decisions automatically. And finally, the HLS metaprogram edits C++ kernels (AST-driven insertion/tuning, data-type/reuse updates) and applies QHS with virtual layers, propagating precision constraints across cascaded operations.
@@ 4.8. Collaborative Optimization Mechanics
LSTM (Que et al., 2020b) MNIST Sequence classification
@@ 5.1. Experimental Setup
[…]
Although our quantitative evaluation uses Jet-DNN, VGG-7, ResNet-9, and LSTM, MetaML-Pro’s compute is governed by its controller and task structure rather than these particular networks. Concretely, the outer Bayesian Optimization (BO) loop runs with a fixed iteration budget (the number of iteration and orderings by default) and convergence criteria that are agnostic to network depth/width (Algorithm. 1), while each inner $O$-task scales gently with model size. Parallel FORK/REDUCE exploration and a thread-pool scheduler further amortize cost across orderings, and the meta-model caches intermediate artifacts across stages to avoid recomputation. These properties allow MetaML-Pro to scale to larger CNN and RNN models. We aim to extend our approach to emerging Transformer families in future work by plugging in the appropriate $\lambda$-tasks and kernels.
@@ 5.2. Single $O$-task Strategies
Figure 10. The auto-pruning algorithm applied to models with binary search direction shown. Omitting step s1 for visibility. The blue arrow indicates an accuracy loss >> user threshold; red denotes the optimal pruning rate.
@@
Figure 12. (a) Quantized bitwidth of each layer in VGG7. (b) Quantized bitwidth of each layer in JetDNN.
@@
The design flow for the pruning strategy is illustrated in Fig. 9(a) and the search steps and direction are shown in Fig. 10. Both the tolerance ($\leq\alpha_{p}$) and threshold ($\beta_{p}$) values are initially set to 2% in this strategy. The effectiveness of the auto-pruning algorithm is demonstrated in Fig. 11. Fig. 11 (a) and (c) depict the pruning rate and accuracy for Jet-DNN and ResNet9 in each step, while (b) and (d) show the resources utilization on Zynq 7020 and U250. As the pruning rate increases, hardware resource requirements, particularly DSPs and LUTs, decrease, leading to improved FPGA performance. The design candidate with the highest pruning rate within the allowed tolerance is selected. In addition, Fig. 10(c)(d) shows that auto-pruning steadily reduces DSP/LUT utilization on U250 for ResNet-9 while maintaining accuracy within the configured $\alpha_{p}$, mirroring the Jet-DNN trend and indicating robustness across CNN families.
@@
This section showcases the evaluation results of the quantization (Q) optimization applied to multiple DNN models within the hardware (HLS) optimization space. The design flow for the quantization strategy is illustrated in Fig. 9(b).
Fig. 12 shows the precision of the weights, biases and output of each virtual layer of the VGG7 model after being tuned by the quantization strategy with $\alpha_{q}$ set to 1%. Table 3 shows how the quantization affected key evaluation metrics relating to the performance and resource usage of 2 DNN designs. With $\alpha_{q}$ set to 1%, the proposed QHS quantization algorithm reduces DSP usage by a factor of 4.9 for the VGG7 model, and 8.5 for the JetDNN model. When $\alpha_{q}$ is increased to 5%, the designs are further compacted, but with a larger real accuracy loss. This table highlights how varying $\alpha_{q}$ levels affect model accuracy, DSP, LUT, and FF usage, illustrating the trade-offs between resource savings and accuracy as well as the effectiveness of the proposed QHS quantization algorithm.
@@ 5.3. Multiple $O$-task Strategies
With our framework, new strategies can be derived by building and revising a design-flow architecture. For instance, by inserting a scaling $O$-task before the pruning $O$-task in Fig. 9(a), a custom combined strategy, which conducts scaling-then-pruning (scaling $\rightarrow$ pruning), can be created with results shown in Fig. 11(e) and (f). The new optimal pruning rate is 84.4%, lower than the previous 93.8% (Fig. 11(a)), due to reduced redundancy from the preceding scaling task. By switching the order of the $O$-tasks, a different optimization strategy performing pruning-then-scaling (pruning $\rightarrow$ scaling) is achieved, resulting in a 0.7% accuracy drop after one scaling step, as seen in Fig. 11(g). Moreover, the three optimization $O$-tasks, pruning, scaling, and quantization, can be integrated into a single automated cross-stage strategy to enhance both performance and hardware efficiency, as illustrated in Fig. 9(c). We discuss the effects of different design-flow architectures with various combinations and orders in Section 5.7, and various tolerable loss in Section 5.8.
@@ 5.4. Branching Flow
Fig. 14 presents the results of applying this design flow to an LSTM model on the MNIST dataset. Specifically, Fig. 14(a) illustrates the pruning rate and accuracy at each step using the PRUNING $O$-task. The tolerance is set to less than $\alpha_{p}$ (2%) in this design. Figs 14(b) and (c) show the resource utilization of the LSTM design after each pruning step respectively on an AMD KU115 FPGA and on an Intel A10 1150 FPGA. The DSP consumption is reduced from 6011 (108%) to 2101 (38.1%) on the AMD FPGA after the final pruning rate is optimized to be 71.9%. Compared to AMD’s HLS compiler, which prefers DSP blocks, Intel’s HLS compiler tends to favor the use of soft multipliers for implementation. As shown in Fig. 14(c), most of the computation kernels are implemented using logic resources rather than DSP blocks. It illustrates that the O-task ports across AMD/Intel with vendor-aware resource trade-offs.
While this design flow currently supports two types of FPGAs, it can be extended to include additional paths such as ASIC technologies. We leave the support for ASIC targets as our future work. Moreover, this evaluation underscores the flexibility of our approach in utilizing the same software optimization task, specifically PRUNING, across multiple hardware targets.
@@ 5.6. Bottom-up Flow
Fig. 16 reveal two distinct flows in the optimization strategies: top-down, where information flows from the DNN to the hardware stage, and bottom-up, the reverse. We have automated both flows by customizing the $K$-task BRANCH with a user-defined predicate function to activate or stop the bottom-up flow if the resulting design overmaps. The BRANCH task also supports a user-supplied action function that is triggered when the predicate condition is true. For our current strategies, the action function changes the CFG section of the meta-model, increasing the accuracy tolerance parameters $\alpha_{p}$, $\alpha_{s}$, and $\alpha_{q}$ (if applicable) for the next iteration. Note that our current strategies in Fig. 16 have two DSE loops: an inner-loop that is codified within each $O$-task and an outer-loop supported by the bottom-up flow. Users can develop more complex strategies by customizing the outer loop, for instance, by updating the bottom-up condition and parameter tuning.
@@ 5.7. Combination and Order
The search for optimal combinations and sequences with multiple $O$-tasks can be computationally demanding due to the exponential growth of the search space with the number of $O$-tasks. Advanced search methods can replace the current brute-force search. Recent studies (Kurek et al., 2016; Ferretti et al., 2022; Wu et al., 2021) demonstrate surrogate models can expedite the search. Our framework facilitates the deployment of such techniques, enabling users to interface a search algorithm with the hardware design with minimal programming effort.